Xavier works in an amusement park and is helping decorate it with strands of lights. This morning, he used a total of 28 strands of lights to decorate 4 bushes and 1 tree. This afternoon, he strung lights on 3 bushes and 1 tree, using a total of 24 strands. Assuming that all bushes are decorated one way and all trees are decorated another, how many strands did Xavier use on each
Answer by ikleyn(52775)

.
Xavier works in an amusement park and is helping decorate it with strands of lights.
This morning, he used a total of 28 strands of lights to decorate 4 bushes and 1 tree.
This afternoon, he strung lights on 3 bushes and 1 tree, using a total of 24 strands.
Assuming that all bushes are decorated one way and all trees are decorated another,
how many strands did Xavier use on each
~~~~~~~~~~~~~~~~


        This problem ideally suits to solve it MENTALLY.


As you read the problem, you should notice in your mind, that the difference 
between work done in the morning and in the afternoon is 4-3 = 1 bush and 28-24=4 strands of lights.


So, you conclude that Xavier uses 4 strands of lights to decorate each bush.


Thus this morning he used 4*4 = 16 strands of lights to decorate 4 bushes,
and the rest 28-16 = 12 strands of lights to decorate one tree.


ANSWER.  4 strands of lights for each bush and 12 strands of lights for each tree.


CHECK the number of strands of lights used in the afternoon:  4*3+12 = 24.   ! correct !

Solved.